I need help with this finance retirement question!

Agnes is 40 years old. She wants to know how much she should be saving each year for retirement. Below are the specifics:

  • She wants to retire at 60 and expects to live until she’s 90

  • She currently makes $45,000 and expects that to increase each year with inflation (2%). She thinks she will need about 70% of that to live on in retirement.

  • She has $40,000 in an RRSP

  • Her investments are earning a real rate of 7%

  • When she retires she will move her investments into a more conservative portfolio and earn 3% per year.

  • She expects that CPP will be about $15,000 per year.


1. How much will she have to save each year to make up the difference? (Note you will need to have completed the first two parts of this problem to do this one.)


2. Given this information, how much will she have saved when she is 60?


3. How much does she need to have saved at the time she retires?
Answer #1

G11 X =PV(G7, 68,-G9,,1) A B с D E F H 1 1 2 3 4 (66868*0.7) 5 6 Number of years to retire =60-40 Inflation rate=2% Earning iG18 =PMT(G14,615,616,-G17) A B с D E H | 13 14 15 16 Rate Nper Pv FV PMT TOTAL SAVINGS REQUIRED PER YEAR Nominal Return beforG26 X =FV(G22, G23,-G24,-G25) А B с D E F G H 1 13 14 Rate Nper 15 16 Py 9.00% 20 $40,000 $944,968 $14,089 17 FV PMT 18 19 20